Write the equation of the line passing through (4, 4) and (4, 2).
A) y = 4 B) y = –2x C) x = 4 D) y = x + 4

Answer by hailu(13)

Determine what the slope is 1st.
m=(2-4)/(4-4)=-2
Take any 1 of the 2 ordered pairs
Lets use (4,2)
Given: m= -2, (4,2)
Linear equation:
y=mx+b
using this equation,substitute the variables with numbers given.
Solve for (b)

y=2,x=4,m=-2
2=-2(4)+b
2=-8+b
2+8=-8+8+b
10=b
So now you solved for m and b, then substitute into the linear equation. But u never substitute for x and y for your answer.
Y=mx+b
Y=-2x+10
i know it's not listed in ur choices but this is the answer i got